# Optical-Cavity Manipulation Strategies of Conical Intersections Mediated Singlet Fission Systems

### v1

@@ Abstract
We offer a theoretical perspective on simulation and engineering of polaritonic conical-intersection-driven singlet-fission (SF) materials. Using rubrene as an example and applying the numerically accurate Davydov-Ansatz methodology, we derive dynamic and spectroscopic responses of the system and demonstrate key mechanisms capable of SF manipulation, viz. cavity-induced enhancement/weakening/suppression of SF, population localization on the singlet state via engineering of the cavity-mode excitation, polaron/polariton decoupling, collective enhancement of SF. We outline unsolved problems and challenges in the field, and share our views on the development of the future lines of research. We emphasize the significance of careful modeling of cascades of polaritonic conical intersections in high excitation manifolds and envisage that collective geometric phase effects may remarkably affect the SF dynamics and yield. We argue that microscopic interpretation of the main regulatory mechanisms of the polaritonic conical-intersection-driven SF can substantially deepen our understanding of this process, thereby providing novel ideas and solutions for improving conversion efficiency in photovoltaics.
With the constant emergence of novel ideas, concepts, and breakthroughs, the boundaries between disciplines have become increasingly blurred, leading to the ascent of highly interdisciplinary fields. One such example is the intersection of molecular physics/chemistry, photonics, and materials science, which gave a birth to optical cavity molecular science and optical cavity materials science Ribeiroa ; Vidal21a ; Bhuyan . By harnessing the strong coupling between photonic modes of an optical cavity (or localized surface plasmon modes of nanoscale structures) and electronic/vibrational degrees of freedom (DOFs) of molecular systems, hybrid electron-vibration-photonic states can be formed Vidal22 ; Huo ; Uji23 ; Rubio23 and characterized by means of nonlinear ultrafast spectroscopy Fassioli ; Volodya23 .
Quantum electrodynamics (QED), the theoretical tool indispensable for modeling/understanding cavity-induced phenomena, was developed in the 1948-1949s Schwinger ; Feynman ; Tomonaga ; Dyson and recently heralded its 75th anniversary (see Orzel for a mini review). The polariton line was observed in the ZnO crystal in 1965 Thomas65 . The light-matter strong-coupling regime was demonstrated on phonons in the thin LiF film in 1975 Zhizhin75 and realized for a single Rydberg atom in the superconducting cavity in 1985 Mueller85 ; Klein87 . In the early 1990s, the strong coupling between excitons and photonic modes was experimentally demonstrated in semiconductor quantum dots Weisbuch . Since then, formation of polaritons owing to the strong coupling between optical-cavity/surface-plasmon modes and molecular excitons in various systems and under different scenarios have attracted significant attention, and a novel area of research known as polaritonic photophysics/photochemistry (PPP) has been established Ribeiroa ; Vidal21a ; Bhuyan ; Vidal22 ; Huo ; Uji23 ; Rubio23 ; Fassioli ; Volodya23 .
@@
PPP deals with a large variety of processes, for example photoisomerization Fregoni ; Fregoni1 , photoinduced reactions in optical cavities or on metal substrates Yang ; Lather , electron and energy transfer processes Mandal ; Campos . Several intriguing phenomena have been discovered in the PPP studies. For instance, cavity-induced modification of the excitonic transport occurs mainly through the delocalized polariton modes rather than tunneling Schachenmayer . Cavities may enhance the exciton coherence length, as well as alter and even eliminate vibronic coupling in J-aggregates Spano15 ; Spano16 . Organic semiconductors showcase increasing conductivity in optical cavities Orgiu , and Raman scattering is enhanced owing to the presence of polaritons Mueller . Strong light-matter coupling has been experimentally demonstrated for chromosomes of photosynthetic bacteria Coles , photosynthetic complexes Tsargorodska , and even for living Chlorobaculum tepidum bacteria InVivo . These and other phenomena demonstrate that PPP holds potential for exciting practical applications, e. g., for engineering novel quantum materials CavityMaterials and development of polaritonic devices KC16 .
One of such applications, which holds a great promise for artificial light harvesting Artificial22 ; Artificial22a and can be substantially improved by the PPP machinery, is singlet fission (SF). It is a spin-allowed photophysical process discovered over fifty years ago SF1 , wherein a molecular material absorbs a visible-light photon generating a singlet exciton S which subsequently undergoes a spin-allowed transformation into two triplet excitons TT Smith1 ; Smith2 ; Rev1 ; Rev2 ; JCP_SF ; Rev3 . Nearly a decade ago, it was shown that dynamics of excitons can be altered drastically by embedding them in an optical cavity Schachenmayer . The idea of utilizing cavities to enhance SF was later proposed, drawing inspiration from the collective enhancement of SF efficiency observed in acene aggregates Mart . Then the pivotal role of the formally dark states for the amplification of SF in the strong-coupling regime in organic materials was demonstrated CavitySF .

### v2

# Optical-cavity manipulation strategies of singlet fission systems mediated by conical intersections: insights from fully quantum simulations

@@ Abstract
We offer a theoretical perspective on simulation and engineering of polaritonic conical-intersection-driven singlet-fission (SF) materials. We begin by examining fundamental models, including Tavis-Cummings and Holstein-Tavis-Cummings Hamiltonians, exploring how disorder, non-Hermitian effects, and finite temperature conditions impact their dynamics, setting the stage for studying conical intersections and their crucial role in SF. Using rubrene as an example and applying the numerically accurate Davydov-Ansatz methodology, we derive dynamic and spectroscopic responses of the system and demonstrate key mechanisms capable of SF manipulation, viz. cavity-induced enhancement/weakening/suppression of SF, population localization on the singlet state via engineering of the cavity-mode excitation, polaron/polariton decoupling, collective enhancement of SF. We outline unsolved problems and challenges in the field, and share our views on the development of the future lines of research. We emphasize the significance of careful modeling of cascades of polaritonic conical intersections in high excitation manifolds and envisage that collective geometric phase effects may remarkably affect the SF dynamics and yield. We argue that microscopic interpretation of the main regulatory mechanisms of polaritonic conical-intersection-driven SF can substantially deepen our understanding of this process, thereby providing novel ideas and solutions for improving conversion efficiency in photovoltaics.
I INTRODUCTION
Cavity molecular science is a rapidly developing interdisciplinary field that has gained significant attention in recent yearsGalego ; Ribeiro2018 ; Mandal . It explores the interaction between molecules and confined electromagnetic fields, such as those within optical or plasmonic cavities. By leveraging the strong coupling between light and matter and the highly hybridized electron-photon-phonon states thus formed, this field aims to manipulate chemical reactions, energy transfer processes, and molecular dynamics in novel manners. Applications have been found in diverse scenarios such as photo isomerization, singlet fission (SF), catalysis enhancement, selective bond manipulation, and intracavity electron transfer processes Orgiu ; Shalabney1 ; Riedinger ; Bienfait ; Vendrell . A typical example is the mixed light-matter quasi particles formed when the interaction between excitons and photons is strong enough, known as polaritons. In 1997, Agranovich et al. proposed a theory of strong light-matter coupling in organic molecules Agranovich , and in 1998, Lidzey et al. conducted the first experimental verification of this theory LIDZEY1 ; LIDZEY2 . In 2011, Schwartz et al. Schwartz introduced a photochemical coupling mechanism for optical switches under strong light, marking the beginning of advancements in polariton optics. Recent research has revealed that the strong coupling between molecular vibrations and a cavity mode can modify the rate constant of ground-state reactions, even in the absence of external driving. A notable characteristic observed in these experiments is the alteration of the vibrational spectrum, attributed to the formation of molecular vibrational polaritons – a hybrid state arising from the interaction between molecular cavity modes and vibrational modes. The existence of vibrational polaritons and changes in chemical reactivity can be characterized by strong resonance Kowalewski23 . Thus, controlling light-matter coupling can be leveraged for polariton transport engineering and the optimization of various photo-induced reactions Vendrell ; JYZ23a .
@@ I INTRODUCTION
Six decades ago, Jaynes and Cummings drew inspiration from studies on quantum coherence in nuclear magnetic resonance and the solution to the Schrödinger equation for two-level systems in a rotating magnetic field to explain deviations between Fermi’s Golden Rule and perturbation theory in microwave resonators. This aligned with Rabi’s findings in 1937. As a result, the renowned Jaynes-Cummings (JC) model JAYNES was developed, describing the interaction between a single two-level atom and a single-mode cavity field under the rotating-wave approximation (RWA). Despite its simplicity, the JC model offers an elegant framework for understanding light-matter interactions. In recent years, various experimental setups have been devised to implement such a system BIRNBAUM ; MAUNZ ; WALLQUIST . Further research has revealed that the JC model exhibits significant complexity in terms of nonlinearity, leading to the development of various extended JC models. Among them, the most noteworthy are the Tavis Cummings (TC) model TAVIS and its variant without RWA, namely, the Dicke model DICKE . One of the most representative features of these Hamiltonians is a second-order quantum phase transition (QPT) from the normal phase to the super-radiant phase in the thermodynamic limit BAUMANN ; KLINDER ; BADEN , a collective phenomenon characterized by condensation of a macroscopic number of photons or atoms.
The interaction between the system and vibrational degrees of freedom (DOFs) can influence molecular behavior within the optical cavity, necessitating the incorporation of phonon modes into the TC model to create a composite framework. This composite model, known as the Holstein-Tavis-Cummings (HTC) model, is widely utilized to study various aspects of polaritonic transport. By tuning the interaction between atoms and cavity fields, steady-state exciton transport can be achieved, and factors influencing atomic transitions, as well as photon emission and absorption processes, can be examined. For the more complex HTC model, approximate analytical and numerical methods are typically employed to investigate its dynamic properties LIU ; Takahashi ; WU ; Spano15 ; Herrera2016 ; HERRERA2 ; HERRERA3 ; Cwik ; HERRERA4 .
Strong light-matter coupling has been experimentally demonstrated for chromosomes of photosynthetic bacteria Coles , photosynthetic complexes Tsargorodska , and even for living Chlorobaculum tepidum bacteria InVivo . These and other phenomena demonstrate that polaritonic photophysics and photochemistry (PPP) holds potential for exciting practical applications, e.g., for engineering novel quantum materials CavityMaterials and development of polaritonic devices KC16 . One of such applications, which holds a great promise for artificial light harvesting Artificial22 ; Artificial22a and can be substantially improved by the PPP machinery, is SF, a spin-allowed photophysical process discovered over fifty years ago in which a single photoexcited singlet exciton is converted into two triplet excitons (TT) SF1 ; Smith2010 ; Smith2 ; Rev1 ; Rev2 ; JCP_SF ; Rev3 ; Musser ; HGD20 ; Miyata . By effectively doubling the number of generated excitons from one photon, SF offers a pathway to exceed the Shockley-Queisser limit and thus enhance power conversion efficiency in solar cells Rao2017 ; Smith2010 ; Congreve2013 . As an all-organic and cost-effective approach to broaden spectral utilization, SF has stimulated extensive research aimed at understanding its underlying principles, optimizing its yields, and integrating it into diverse molecular architectures and solid-state platforms Berkelbach2013 ; Margulies2016 . Nonetheless, controlling the rate and coherence of SF remains challenging, partly due to the intricate interplay of electronic interactions, vibronic couplings, and complex many-body dynamics Monahan2015 ; Margulies2016 .